# Booster: Tackling Harmful Fine-tuning for Large Language Models via Attenuating Harmful Perturbation

### v3

@@ 5.1 Setup
Datasets and models. For the alignment task, we use the alignment dataset and harmful dataset from (Rosati et al., 2024b), which is enriched from BeaverTails (Ji et al., 2023). In the alignment stage, we sample 5000 instances to construct the alignment dataset, and another 5000 instances to construct the harmful dataset. The data in harmful dataset are in the same distribution but are different from those harmful data mixed in the user dataset. For fine-tuning task, we consider SST2(Socher et al., 2013), AGNEWS(Zhang et al., 2015), GSM8K(Cobbe et al., 2021) and AlpacaEval (Li et al., 2023) as the user fine-tuning task. To simulate the harmful fine-tuning attack, we mix $p$ (percentage) of unsafe data from BeaverTail with $1-p$ benign fine-tuning data over a total number of $n$ samples. We use Llama2-7B (Touvron et al., 2023), and two SOTA architecture Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024) for evaluation. In our experiment, the default setting is $p=0.1$ and $n=1000$ (specially, $n=700$ for AlpacaEval) unless otherwise specified. All the experiments are done with an H100-80G.
Metrics. We follow Huang et al. (2024c; b) to use two metrics for evaluation of model’s performance.
  • Finetune Accuracy (FA). The accuracy of the testing dataset from the corresponding finetune task. We give a detail of how to measure the accuracy for different tasks in Appendix A.
@@
To calculate harmful score, we sample 1000 instructions from the testing set of BeaverTails (Ji et al., 2023). To obtain finetune accuracy, we sample 872, 1000, 1000, and 122 samples respectively from finetuning dataset SST2, AGNEWS, GSM8K and AlpacaEval. Both the two metrics are measured on the fine-tuned model (after two-stage training).
Baselines. We use four baselines for comparison. SFT is the vanilla supervised fine-tuning solution. Lisa (Huang et al., 2024b) is a fine-tuning stage solution, and Vaccine(Huang et al., 2024c) and RepNoise (Rosati et al., 2024a) are two alignment stage solutions for the harmful fine-tuning issue.
Training details. We follow (Huang et al., 2024c; b; Hsu et al., 2024) to utilize LoRA (Hu et al., 2021) for efficient LLM training. The rank of the adaptor is set to 32, and the LoRA’s alpha is 4. For alignment, we use AdamW as optimizer (Loshchilov et al., 2017) with a learning rate 5e-4 and a weight decay factor of 0.1. For fine-tuning tasks, we use the same optimizer with a smaller learning rate 1e-5 following (Huang et al., 2024c). We train 20 epochs for alignment, and 20 epochs for fine-tuning with SST2, AGNEWS and GSM8K, and 100 epochs for AlpacaEval. For both alignment and fine-tuning stage, we use the same batch size of 10. See Appendix A for details.
@@ 5.2 Main Experiments
Robustness to harmful ratio. We first show in Table 1 how the harmful ratio affects the model’s safety. As shown, compared to SFT, Booster in average achieves 22.64% of lower harmful score, and 2.64% higher finetune accuracy on the downstream. Particularly, we observe that Booster achieves significantly higher finetune accuracy compared to SFT when p=0 (i.e., clean). We conjecture that the reason is that the alignment with SFT hurts the finetune performance due to over-fitting (i.e., the model learns to use refusal answer even not for harmful prompt). While Booster does not specifically target on solving the over-fitting problem, the attenuating regularizer avoids the model to minimize one objective, i.e., alignment loss. With all the harmful ratio, we see that Booster maintain consistently better harmful score and finetune accuracy compared to baseline, but we do observe that the harmful score is rising with more percentage of harmful data, which unfortunately is the common weakness of the alignment-stage solutions.
@@
Robustness to fine-tuning sample number. Next we show in Table 2 how different fine-tuning sample number affects the defense performance. As shown, Booster in average achieves 28.76% reduction of harmful score compared to SFT, with 2.13% increase of finetune accuracy. When the fine-tuning sample is small, Booster achieves significantly higher finetune accuracy than SFT. Analogy to our analysis above. the reason is that with less fine-tuning samples, the overfitting over the alignment data may not be well mitigated by SFT. However, Booster can mitigate the overfitting of alignemnt data with the added attenuating regularizer.
Table 2: Performance analysis for different sample number.
@@
Generalization to fine-tuning datasets. Table 3 shows the comparison results over four different fine-tuning datasets. Results show that in respectively achieves 25.4%, 23.6%, 8.4%, and 4.0% of harmful score reduction compared to SFT, and achives the highest average finetune accuracy among all the baselines. The performance is significantly improved compared to two other alignment stage solutions RepNoise and Vaccine, with respectively 19.68% and 16.7% reduction of harmful score. The experiment justifies that the proposed method can be extended to more complicated fine-tuning tasks, e.g., GSM8K and AlpacaEval.
@@
Generalization to models. The above experiment is done with Llama2-7B. In Table 4 we show that the proposed method can be extended to two latest SOTA model architectures, i.e., Gemma2-9B, and Qwen2-7B. In average, Booster achieves 34.14% reduction of harmful score and 0.71% improvement of finetune accuracy. Particularly, Booster maintains an astounding 1.6% harmful score with 95.64% of finetune accuracy for Qwen2-7B, which justifies its generalization to SOTA LLM architecture trained on massive data/token scale.
@@ 5.3 Statistical/System Analysis
Statistical Analysis. Figure 3 shows the statistical comparison results between SFT and Booster. We derive the following observation of the three statsitics.
  • Harmful Score. As shown in the left Figure, the harmful score of Booster is not significantly improved with training epochs while SFT without defense is rising to a high value and seems to have trend to grow even after 2000 steps.
  • Harmful Training Loss. The middle figure demonstrate that SFT initially has a high harmful training loss, but it is drastically reduced with the fine-tuning epochs. On contrary, the aligned model of Booster initially has a relatively low harmful training loss, but it decreases slowly with the fine-tuning epochs. After 2000 steps, the harmful training loss of SFT is significantly lower than that of Booster. The result of Booster coincide our design motivation to attenuate the harmful loss reduction rate.
  • Harmful Testing Loss. The right figure show the change of harmful testing loss (i.e., loss over harmful data unseen on the fine-tuning stage). The trend of this statistic is mostly the same with harmful training loss. This is in expectation because the harmful testing data is in the same data distribution with the harmful training data, and they model fit in the harmful training data can be generalized to the unseen harmful data.
System Evaluation. Table 5 shows the system evaluation results. Results indicate that the superior performance of Booster, unfortunately, does not come for free. The Clock time of Booster used for alignment is approximately three times of SFT. This is because for each optimization step, Booster needs to evaluate the gradient for three times, which triple the training time. Booster in the alignment stage also incurs approximately additional 8.53GB memory compared to SFT. The overhead comes from the storage of the the three gradient vectors, as well as the storage of one batch of harmful data. However, Booster is still superior to another alignment stage solution RepNoise, which requires 0.83h more clock time and 14.61GB more GPU memory. Additionally, we need to note that Booster, as well as two other alignment stage solutions do not incur additional computation in the fine-tuning stage. This is particularly important, because alignment only needs to be done once, and the aligned model can serve as the foundation model for millions of fine-tuning requests. This means that the overhead is only a one-time cost, but a fine-tuning stage solution,e.g., Lisa needs to incur overhead for every fine-tuning request. In summary, our claim is that Booster requires more computation/memory overhead, but this is acceptable due to its nature as an alignment-stage solution.
Table 5: System evaluation for different methods. Booster introduces extra overhead in alignment stage because it needs extra forward/backward passes.
@@ 5.4 Hyper-parameter Analysis
Impact of attenuating regularizer’s intensity $\lambda$. Table 6 shows how $\lambda$ affects the defense performance. As shown, setting $\lambda=0$ in Booster reduces the solution to SFT with high harmful score. When $\lambda$ is too large, defense performance downgrades with increased harmful score. This probably is because the model cannot optimize well with the safety loss with too large $\lambda$. Therefore, $\lambda$ needs to be carefully tune to the right value to guarantee practical performance of Booster.
@@
Impact of inner step size $\alpha$. Table 7 show how the inner step size affects Booster’s defense performance. The inner step aims to take one normalized step towards the harmful gradient direction to simulate the harmful perturbation, and its step size $\alpha$ should be carefully tuned to guarantee practical performance. As shown in the table, when $\alpha=0$, the defense will reduce to SFT solution, and therefore the model’s harmful score cannot be properly reduced. When $\alpha$ is set to be too large, i) the model optimization might incur instability (e.g., when $\alpha=0.5$), and ii) the defense will be ineffective. This probably is because that a too large step size cannot simulate the harmful perturbation, and therefore Booster with the attenuating regularizer will not be useful when encountering the real harmful perturbation.
@@
Impact of number of harmful samples. In Booster, we utilize a harmful dataset to simulate the harmful perturbation, and with which we immunize the model. In table 8, we show how the number of harmful samples being used affects the performance of Booster. As shown, a small amount of harmful samples, e.g., 50 instances is enough for an accurate simulation of harmful perturbation, and thereby ensuring a low harmful score. However, as expected, a too small amount of harmful instances used in alignment, e.g., 5 instances, is not sufficient to approximate the harmful perturbation, and the defense fails in this setting.
@@ 5.5 Alternative Design
Vaccine+Booster. Booster can be combined with existing alignment-stage solution Vaccine (Huang et al., 2024c) for performance enhancement. For space limitation, we discuss our detailed implementation of Vaccine+Booster in Appendix D. Table 9 shows the performance of Vaccine+Booster compared with the their original implementations. As shown, Vaccine+Booster in average achieves 3.88% reduction of harmful score with 0.74% drop of finetune accuracy compared to the original Booster. However, it is notable that Vaccine+Booster may need more computation because it requires one more forward/backward pass of the model with safety alignment data.
Table 9: Performance of Vaccine+Booster. The number of alignment data used in the alignment stage is 500.
@@ 5.6 Visualization
As follows, We show how different methods respond to the malicious prompt. We used the finetuned model over default attack setting for evaluation. As shown, Booster is able to give refusal answer to the sensitive question while other methods cannot.
@@ 6 Conclusion
In this paper, we first show that harmful perturbation in the fine-tuning stage leads to the decrease of harmful loss and therefore leading to the break of alignment. To mitigate this issue, we propose to solve a perturbation minimization problem by constraining the gap of harmful loss before and after a simulated perturbation. The proposed problems are solved by a iterative gradient method named Booster. Despite its embarrassing simplicity, experimental results show that the proposed method is able to mitigate the risk of fine-tuning while maintaining the finetune accuracy.
Appendix A Experimental Details
@@
Hyper-parameters. For alignment stage, the learning rate is 5e-4 with batch size 10. We train 20 epochs for 5000 pieces of alignment data. The alignment data is sampled from BeaverTails Ji et al. (2023). For fine-tuning stage, the learning rate is 1e-5 with batch size 10. We also train 20 epochs for 1000 (default) fine-tuning samples. The fine-tuning samples are a mixture of benign fine-tuning samples and harmful instances. The harmful instances are sampled from BeaverTails Ji et al. (2023), while the benign fine-tuning samples are sampled from dataset for specific task. For example, we sample benign fine-tuning samples for GSM8K training set [^3] https://huggingface.co/datasets/openai/gsm8k for GSM8K task. By default, we adopt harmful ratio $p=0.1$ and $n=1000$ as the ratio of harmful data and the number of the total fine-tuning samples. Specially, for AlpcaEval, the number of fine-tuning samples is 700 because the lack of training data for this task.
@@
We follow Huang et al. (2024c; b; a) to construct the prompt templates of different tasks. For SST2, the instruction is "Analyze the sentiment of the input, and respond only positive or negative", the input is the according sentence in SST2 dataset, and the output is the real sentiment label in SST2. In testing time, we ask the model to complete the output given instruction and input. The answer is classified to be correct if it matches the label. For GSM8K, the instruction is the real mathematics qeustion in GSM8K, and the output is the real answer. In testing time, the answer is considered to be correct if the final answer is able to given by the model. For AlpacaEval, we use the GPT4’s high quality instruction/answer pair as the demonstration data. In testing time, we evaluate the helpfulness of the output of the model over unseen general prompt via calling chatGPT’s API.
Appendix B Baseline Descriptions
We in this section first briefly introduce the how the existing baselines are implemented in our experiments.

### v4

@@ 5.1 Setup
Datasets. For the alignment task, we use the alignment dataset and harmful dataset from (Rosati et al., 2024c), which is enriched from BeaverTails (Ji et al., 2023). In the alignment stage, we sample 5000 instances to construct the alignment dataset, and another 5000 instances to construct the harmful dataset. The data in harmful dataset are in the same distribution but are different from those harmful data mixed in the user dataset. For fine-tuning task, we consider SST2(Socher et al., 2013), AGNEWS(Zhang et al., 2015), GSM8K(Cobbe et al., 2021) and AlpacaEval (Li et al., 2023a) as the user fine-tuning task. To simulate the harmful fine-tuning attack, we mix $p$ (percentage) of unsafe data from BeaverTail with $1-p$ benign fine-tuning data over a total number of $n$ samples.
Models. We use Llama2-7B (Touvron et al., 2023), and two SOTA architecture Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024) for evaluation.
In our experiment, the default setting is $p=0.1$ and $n=1000$ (specially, $n=700$ for AlpacaEval) and Llama2-7B as the base model unless otherwise specified.
Metrics. We follow Huang et al. (2024e; d) to use two metrics for evaluation of performance.
  • Finetune Accuracy (FA). The accuracy of the testing dataset from the corresponding finetune task. We give a detail of how to measure the accuracy for different tasks in Appendix B.
@@
To calculate the harmful score, we sample 1000 instructions from the testing set of BeaverTails (Ji et al., 2023). To obtain finetune accuracy, we sample 872, 1000, 1000, and 122 samples respectively from finetuning dataset SST2, AGNEWS, GSM8K, and AlpacaEval. Both the two metrics are measured on the fine-tuned model (i.e., after two-stage training).
Baselines. We use four baselines for comparison. SFT is the vanilla supervised fine-tuning solution. Lisa (Huang et al., 2024d) is a fine-tuning stage solution, and Vaccine(Huang et al., 2024e) and RepNoise (Rosati et al., 2024b) are two alignment stage solutions for the harmful fine-tuning issue.
Training details. We follow (Huang et al., 2024e; d; Hsu et al., 2024) to utilize LoRA (Hu et al., 2021) for efficient LLM training. The rank of the adaptor is set to 32, and the LoRA’s alpha is 4. For alignment, we use AdamW as optimizer (Loshchilov et al., 2017) with a learning rate 5e-4 and a weight decay factor of 0.1. For fine-tuning tasks, we use the same optimizer with a smaller learning rate 1e-5 following (Huang et al., 2024e). We train 20 epochs for alignment, and 20 epochs for fine-tuning with SST2, AGNEWS and GSM8K, and 100 epochs for AlpacaEval. For both alignment and fine-tuning stage, we use the same batch size of 10. The default hyper-parameters for Booster are $\lambda=5$ and $\alpha=0.1$ and we use 5000 pieces of harmful data in alignment to simulate harmful perturbation. See Appendix B for details.
@@ 5.2 Main Experiments
Robustness to harmful ratio. We first show in Table 1 how the harmful ratio affects the model’s safety. As shown, compared to SFT, Booster in average achieves 22.64% of lower harmful score, and 2.64% higher finetune accuracy on the downstream task. Particularly, we observe that Booster achieves significantly higher finetune accuracy compared to SFT when p=0 (i.e., clean). We conjecture the reason is that the alignment with SFT hurts the finetune performance due to over-fitting (i.e., the model learns to use refusal answers even are prompted with harmless questions). While Booster does not specifically target on solving the over-fitting problem, the attenuating regularizer avoids the model to minimize one objective, i.e., alignment loss. With all the harmful ratios, we see that Booster maintains consistently better harmful score and finetune accuracy compared to baseline, but we do observe that the harmful score is rising with more percentage of harmful data, which unfortunately is the common weakness of the alignment-stage solutions.
@@
Robustness to fine-tuning sample number. Next, we show in Table 2 how different fine-tuning sample number affects the defense performance. As shown, Booster in average achieves 28.76% reduction of harmful score compared to SFT, with 2.13% increase in finetune accuracy. When the fine-tuning sample is small, Booster achieves significantly higher finetune accuracy than SFT. Analogy to our analysis above, the reason is that with less fine-tuning samples, the overfitting over the alignment data may not be well mitigated by SFT. However, Booster can mitigate the overfitting of alignment data with the added attenuating regularizer.
Table 2: Performance analysis for different sample number for fine-tuning.
@@
Generalization to fine-tuning datasets. Table 3 shows the comparison results over four different fine-tuning datasets. Results show that Booster respectively achieves 25.4%, 23.6%, 8.4%, and 4.0% of harmful score reduction compared to SFT on the four datasets, and achieves the highest average finetune accuracy among all the baselines. The performance is significantly improved compared to two other alignment stage solutions RepNoise and Vaccine, with respectively 19.68% and 16.7% reduction of harmful score. The experiment justifies that the proposed method can be extended to more complicated fine-tuning tasks, e.g., GSM8K and AlpacaEval.
@@
Generalization to models. The above experiment is done with Llama2-7B. In Table 4 we show that the proposed method can be extended to two latest SOTA model architectures, i.e., Gemma2-9B, and Qwen2-7B. On average, Booster achieves 34.14% reduction of harmful score and 0.71% improvement of finetune accuracy. Particularly, Booster maintains an astounding 1.6% harmful score with 95.64% finetune accuracy for Qwen2-7B, which justifies its generalization to SOTA LLM architecture trained on massive data/token scale.
@@ 5.3 Statistical/System Analysis
Statistical Analysis. Figure 3 shows the statistical comparison results between SFT and Booster. We derive the following observation from the three statistics.
  • Harmful Score. As shown in the left figure, the harmful score of Booster is not significantly improved with training epochs while SFT without defense is rising to a high value and seems to have a trend to grow even after 2000 steps.
  • Harmful Training Loss. The middle figure demonstrates that SFT initially has a high harmful training loss, but it is drastically reduced with the fine-tuning epochs. On the contrary, the aligned model of Booster initially has a relatively low harmful training loss, but it decreases slowly with the fine-tuning epochs. After 2000 steps, the harmful training loss of SFT is significantly lower than that of Booster. The result of Booster coincides with our design motivation to attenuate the harmful loss reduction rate.
  • Harmful Testing Loss. The right figure shows the change in harmful testing loss (i.e., loss over harmful data unseen in the fine-tuning stage). The trend of this statistic is mostly the same with harmful training loss. This is in expectation because the harmful testing data is in the same data distribution as the harmful training data, and the model that fits the harmful training data can be generalized to the unseen harmful data.
System Evaluation. Table 5 shows the system evaluation results. Results indicate that the superior performance of Booster, unfortunately, does not come for free. The clock time of Booster used for alignment is approximately three times of SFT. This is because, for each optimization step, Booster needs to evaluate the gradient three times, which triple the training time. Booster in the alignment stage also incurs approximately an additional 8.53GB of memory compared to SFT. The overhead comes from the storage of the three gradient vectors, as well as the storage of one batch of harmful data. However, Booster is still superior to another alignment stage solution RepNoise, which requires 0.83h more clock time and 14.61GB more GPU memory. Additionally, we need to note that Booster, as well as two other alignment stage solutions, do not incur additional computation in the fine-tuning stage. This is particularly important, because alignment only needs to be done once, and the aligned model can serve as the foundation model for millions of fine-tuning requests. This means that the overhead is only a one-time cost, but a fine-tuning stage solution,e.g., Lisa needs to incur overhead for every fine-tuning request. In summary, our claim is that Booster requires more computation/memory overhead, but this is not unacceptable due to its nature as an alignment-stage solution.
Table 5: System evaluation for different methods. Booster introduces extra overhead in the alignment stage because it needs extra forward/backward passes.
@@ 5.4 Hyper-parameter Analysis
Impact of attenuating regularizer’s intensity $\lambda$. Table 6 shows how $\lambda$ affects the defense performance. As shown, setting $\lambda=0$ in Booster reduces the solution to SFT with a high harmful score. When $\lambda$ is too large, defense performance downgrades with the increased harmful score. This probably is because the model cannot optimize well with the safety loss with too large $\lambda$. Therefore, $\lambda$ needs to be carefully tuned to the right value to guarantee the practical performance of Booster.
@@
Impact of inner step size $\alpha$. Table 7 shows how the inner step size affects Booster’s defense performance. The inner step aims to take one normalized step towards the harmful gradient direction to simulate the harmful perturbation, and its step size $\alpha$ should be carefully tuned to guarantee practical performance. As shown in the table, when $\alpha=0$, the defense will reduce to SFT solution, and therefore the model’s harmful score cannot be properly reduced. When $\alpha$ is set to be too large, i) the model optimization might incur instability (e.g., when $\alpha=0.5$), and ii) the defense will be ineffective. This probably is because a too-large step size cannot simulate the harmful perturbation, and therefore Booster with the attenuating regularizer will not be useful when encountering the real harmful perturbation.
@@
Impact of number of harmful samples. In Booster, we utilize a harmful dataset to simulate the harmful perturbation, with which we immunize the model. In table 8, we show how the number of harmful samples being used affects the performance of Booster. As shown, a small amount of harmful samples, e.g., 50 instances is enough for an accurate simulation of harmful perturbation, thereby ensuring a low harmful score. However, as expected, a too small amount of harmful instances used in alignment, e.g., 5 instances, is not sufficient to approximate the harmful perturbation, and the defense fails in this setting.
@@ 5.5 Alternative Design
Vaccine+Booster. Booster can be combined with existing alignment-stage solution Vaccine (Huang et al., 2024e) for performance enhancement. For space limitation, we discuss our detailed implementation of Vaccine+Booster in Appendix F. Table 9 shows the performance of Vaccine+Booster compared with their original implementations. As shown, Vaccine+Booster in average achieves 3.88% reduction of the harmful score with 0.74% drop in finetune accuracy compared to the original Booster. However, it is notable that Vaccine+Booster may need more computation because it requires one more forward/backward pass of the model with safety alignment data.
Table 9: Performance of Vaccine+Booster. The number of alignment data used in alignment is 500.
@@ 5.6 Visualization
In the following, we show how different methods respond to the malicious prompt. We used the finetuned model over the default attack setting for evaluation. As shown in the example (next page), Booster is able to give a refusal answer to sensitive questions while other methods cannot.
@@ 6 Conclusion
In this paper, we first show that harmful perturbation in the fine-tuning stage leads to the decrease of harmful loss and therefore leads to the break of alignment. To mitigate this issue, we propose to solve a perturbation minimization problem by constraining the gap of harmful loss before and after a simulated perturbation. The proposed problems are solved by an iterative gradient method named Booster. Despite its embarrassing simplicity, experimental results show that the proposed method is able to mitigate the risk of fine-tuning while maintaining the finetune accuracy.
7 Ethics statement
We propose in this paper a defense towards harmful fine-tuning attack. The technique itself alone should not pose serious ethical concern. For demonstration purposes, this paper presents a few harmful data that some may find disturbing or offensive, including content that is hateful or violent in nature.
8 Reproducibility Statement
We make the following effort to enhance the reproducibility of our results.
  • For Booster implementation, the pseudo-algorithm Algorithm 1 should be briefly enough to illustrate the algorithm logic. For implementation, we strictly follow the pseudo-algorithm without adding any additional tricks to the code. A link to an anonymous downloadable source is included in our abstract for public verification.
  • To derive the harmful dataset and the alignment dataset, we use the data pair from (Rosati et al., 2024b) with this link: https://huggingface.co/datasets/anonymous4486/repnoise_beavertail. Other used datasets are standard benchmark datasets that can be accessed from Huggingface.
  • We show a brief description of defense baselines in Appendix C.
  • The hyper-parameters and simulation settings are given in Section 5.1. Detailed setting and the hyper-parameter selection logistic can be found in Appendix B.
9 Acknowledgment
This research is partially sponsored by the NSF CISE grants 2302720, 2312758, 2038029, an IBM faculty award, a grant from CISCO Edge AI program. This research is supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA. All the authors truly appreciate the constructive review comments from the anonymous reviewers/ACs during our submissions to ICLR2025.
Appendix A Booster vs. TAR
A concurrent alignment stage solution TAR (Tamirisa et al., 2024), first released on August 1st (approximately one month before Booster’s first release) shares a similar optimization viewpoint to our proposed Booster solution. We then present a simplified idea of TAR (only one inner step in TAR is taken for the sake of easy comparisons with Booster). Notations are also adapted for easy comparison with Booster. Formally, a simplified version of TAR aims to solve the following problem:
where $\tilde{f}(\bm{w})$ is a representation engineering-inspired retain loss for preserving performance on the capabilities proxy dataset, and $h(\bm{w})$ is the entropy loss over harmful data. The second term is a tamper-resistance loss that ensures the negative entropy loss over the harmful data is minimized after one step of a harmful fine-tuning attack. Of note, in original TAR, multiple inner steps are taken to simulate the harmful fine-tuning attack.
In contrast, Booster aims to solve the following problem:
The main difference lies in two main aspects.
  • $f(\bm{w})$ vs. $\tilde{f}(\bm{w})$. For optimizing the alignment loss $f(\bm{w})$, Booster needs to assume the existence of alignment data. This can ensure the model learns to answer harmful prompts in a refusal way. TAR instead optimizes over the retain loss, which requires the existence of a normal proxy dataset.
  • The main difference lies in the second term. TAR aims to directly minimize the negative entropy loss over harmful data after taking the simulated harmful perturbation. In contrast, Booster aims to minimize the harmful loss reduction rate by minimizing $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$. While the difference seems to be subtle, the goals of the two designs, as we mentioned, are different. In our initial attempt, we also experimented on a term similar to tamper resistance loss used by TAR. We minimize over the negative cross-entropy loss but the training would become very unstable with this loss because the model will easily collapse ($h(\bm{w})$ will explode to a large value, and eventually leading the model to constantly repeat a single word for any prompts). Therefore, to guarantee the model’s stability, we eventually adopt the current Booster formulation by minimizing the harmful loss reduction rate.
Appendix B Experimental Details
@@
Hyper-parameters. For the alignment stage, the learning rate is 5e-4 with batch size 10. We train 20 epochs for 5000 pieces of alignment data. The alignment data is sampled from BeaverTails Ji et al. (2023). For the fine-tuning stage, the learning rate is 1e-5 with batch size 10. We also train 20 epochs for 1000 (default) fine-tuning samples. The fine-tuning samples are a mixture of benign fine-tuning samples and harmful instances. The harmful instances are sampled from BeaverTails Ji et al. (2023), while the benign fine-tuning samples are sampled from the dataset for specific tasks. For example, we sample benign fine-tuning samples for GSM8K training set [^3] https://huggingface.co/datasets/openai/gsm8k for GSM8K task. By default, we adopt harmful ratio $p=0.1$ and $n=1000$ as the ratio of harmful data and the number of the total fine-tuning samples. Specially, for AlpcaEval, the number of fine-tuning samples is 700 because of the lack of training data for this task.
@@
We follow Huang et al. (2024e; d; a) to construct the prompt templates of different tasks. For SST2, the instruction is "Analyze the sentiment of the input, and respond only positive or negative", the input is the according sentence in SST2 dataset, and the output is the real sentiment label in SST2. In testing time, we ask the model to complete the output given instruction and input. The answer is classified to be correct if it matches the label. For GSM8K, the instruction is the real mathematics question in GSM8K, and the output is the real answer. In testing time, the answer is considered to be correct if the final answer is able to be given by the model. For AlpacaEval, we use the GPT4’s high-quality instruction/answer pair as the demonstration data. In testing time, we evaluate the helpfulness of the output of the model over unseen general prompt via calling chatGPT’s API.
Appendix C Baseline Descriptions
We in this section first briefly introduce how the existing baselines are implemented in our experiments.